40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 22 01:03:30 2012, Den, Update, PEM, ETMY table X.pdfY.pdfZ.pdf
    Reply  Mon Sep 24 17:03:39 2012, Jenne, Update, PEM, 1X1 rack power 
    Reply  Fri Oct 12 22:44:49 2012, Den, Update, PEM, ETMY table 6x
       Reply  Sun Oct 14 11:41:42 2012, Den, Update, PEM, ETMY table DSC_4739.JPG14OCT2012.pdf14OCT2012.xml
    Reply  Thu Dec 13 01:49:53 2012, rana, Update, PEM, seismometers a.pdf
       Reply  Thu Dec 13 21:21:34 2012, Den, Update, PEM, seismometers gur.pdfgur_fix.pdf
Message ID: 7819     Entry time: Thu Dec 13 01:49:53 2012     In reply to: 7429     Reply to this: 7825
Author: rana 
Type: Update 
Category: PEM 
Subject: seismometers 

Quote:

I've installed Guralp readout box back and it turned out that it does not work with voltage provided from the rack (+13.76 0 -14.94).  +/-12 voltage regulators inside the box convert it to -0.9 0 -12. I've connected the box to +/-15 DC voltage supply to measure seismic motion at the ETMY table. Readout box works fine with +/- 15.

Seismic noise on the ETMY table measured to be a few times higher then on the floor in horizontal direction in the frequency range 50 - 200 Hz. Attached are compared spectrums of X, Y and Z motions.

 Looking at the PEM BLRMS, I noticed that the GUR1Z channel had a much reduced microseism compared to the GUR1X. Looking at the BLRMS screens everything seems ON, although its a mess (too many filters in the banks, etc. - clean this up, PEM people).

 Looking at the Z channel in DTT, I see that the Z spectra looks double high pass filtered below ~1 Hz.     Needs some attention in the daytime.

From Den and Ayaka's elog entry from Nov 29, its clear that this problem is there at that time. It seems that the seismometer was not even hooked up before then. Perhaps Tara returned the seismometer around Thanksgiving and then someone here hooked it up but neglected to log this work? If so, please make an elog now describing the installation of this sensor at the 40m and log any future work which takes place at the 40m lab even if you think it is inconsequential.

Attachment 1: a.pdf  145 kB  | Hide | Hide all
a.pdf
ELOG V3.1.3-